OpenStudy (anonymous):

quadractic formula: x^2-3x=18? will someone help

OpenStudy (anonymous):

You want to get an equation that equals 0, so you subtract 18 for both sides

OpenStudy (anonymous):

x^2-3x-18=18-18 x^2-3x-18=0

OpenStudy (anonymous):

You can factor this, but do you want to use the quadratic formula?

OpenStudy (anonymous):

yes. the quadratic form please

OpenStudy (anonymous):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

We have the quadratic in the form ax^2+bx+c=0

OpenStudy (anonymous):

Can you identify a, b, c?

OpenStudy (anonymous):

a=1 b=3 c=-18?

OpenStudy (anonymous):

a=1, b=-3, c=-18, don't forget your signs

OpenStudy (anonymous):

so now we can plug these numbers into the quadratic formula \[x=\frac{3 \pm \sqrt{(-3)^2-4(1)(-18)}}{2(1)}\]

OpenStudy (anonymous):

why is the 3 positive?

OpenStudy (anonymous):

\[x=\frac{3 \pm \sqrt{9+72}}{2}=\frac{3 \pm \sqrt{81}}{2}=\frac{3 \pm 9}{2}\]

OpenStudy (anonymous):

the 3 is positive because its -b

OpenStudy (anonymous):

-(-3)=3

OpenStudy (anonymous):

ok. i c now.thank you.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

one more question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i have an imaginary number question... when solving the problem my answer is 3i+6/9. Do I reduce both numbers by 3 resulting answer would be i+2/3

OpenStudy (anonymous):

\[\frac{3i+6}{9} or 3i+\frac{6}{9}\]

OpenStudy (anonymous):

1st one or second one?

OpenStudy (anonymous):

1stt

OpenStudy (anonymous):

\[\frac{3(i+2)}{9}\] Yes, you can reduce to \[\frac{i+2}{3}\]

OpenStudy (anonymous):

okay. thanks

OpenStudy (anonymous):

you're welcome